OpenStudy (destinyyyy):

What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1?

OpenStudy (destinyyyy):

I got.. -1=1(3)=B B= -4 m= 1 y=1x-4 y=x-4